Status Report

NASA Space Station On-Orbit Status 12 Aug 2003

By SpaceRef Editor
August 12, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 12 Aug 2003

All ISS systems continue to function nominally, except as noted previously or below.

After wake-up (2:00am EDT), before breakfast and exercise, CDR Yuri Malenchenko completed the periodic Russian MedOps test “Hematokrit” (MO-10), which measures the blood’s red cell count.  [The blood sample was drawn from a finger with a perforator lancet, then centrifuged in two microcapillary tubes in the M-1100 kit’s minicentrifuge, and its hematocrit value was read off the tubes with a magnifying glass.  It is a well-known phenomenon of space flight that red blood cell count (normal range: 30-45%) tends to go down over time.]

Immediately afterwards, Malenchenko configured the Russian payload laptop 3 and performed the MBI-1 SPRUT-K experiment, part of Russian medical research on the distribution and behavior of human body fluids in zero gravity.  [The experiment involves use of  the Sprut (“squid”) securing harness and skin electrodes (cuffs), along with laptop 3, which was then reconfigured to its nominal settings.]

FE/SO Ed Lu completed the fourth 30-day functionality (“health”) test of the HRF GASMAP (Human Research Facility/Gas Analyzer System for Metabolic Analysis Physiology).  Last time done: 7/14.  [After unstowing, cabling and powering up the equipment, he let it run for 3 hrs. in standby mode, then performed a health check and reconfigured the system for a 2-hr. low power rundown.  Afterwards, about 6 hrs. after power-up, the GASMAP was deactivated, disconnected and stowed again.]

Yuri Malenchenko began with the scheduled transfer of liquid waste to the empty Rodnik water storage tanks (BV1, BV2) of 10P/Progress-247.  Today’s operation transferred stored urine from 4 EDV-U containers into tank BV1, from a total of 17 filled EDVs.  The work will be continued tomorrow and Thursday (8/14).  [In preparation for this activity, on 8/4 Yuri had pressurized the folded bladders of the 10P Rodnik water storage tanks.  Each of the two spherical tanks consists of a hard shell with a soft membrane (bladder) composed of elastic fluoroplastic, which is used to expel water from the tank by compressed air pumped into the tank volume surrounding the bladder.]

The crew completed the replacement of the TVIS/treadmill roller bearing assemblies begun yesterday.  [After yesterday’s swap-out of the ten centermost bearings on the left side of the TVIS’ internal truss and removal of the corresponding ten bearings on the right side, Ed and Yuri finished the job today with the installation of the second set of ten roller assemblies.  The crew was congratulated on doing a “fantastic job” on the IFM (in-flight maintenance).]

Ed Lu had a 15-min. time block reserved for stowing any ancillary EMU items that may still have been unstowed after the EMU/spacesuit #3005 resizing and fit checking on 8/7.  [The next EVA/EMU troubleshooting is scheduled for 8/15 (Friday).  Even after Ed’s and the ground’s best attempts to date, EMU #3013 still fails to provide any cooling.  Next step will be for Ed to go underneath the TMG (thermal micrometeoroid garment) outer layer to look for various component failures or clogged filters.  The situation for EMU #3005, which exhibits intermittent or degraded cooling, provides a new signature for consideration, and its data are still being analyzed.  By having Ed Lu perform water flow tests on #3013 on 8/15, component failures could possibly be ruled out that would be unique to having no cooling flow, and converge on a common, more generic issue related to degraded flow or filter clogging of the spacesuits.]

During the time allotted for his TVIS troubleshooting, Ed Lu completed the PFMI (Pore Formation and Mobility Investigation) set-up in the MSG (Microgravity Science Glovebox).  [For this purpose, the MSG was powered up by the ground beforehand, later to be deactivated again.]

The Science Officer also had another go at setting up the EPO (Educational Payload Operations) demonstration of EVA/IVA tools scheduled to be performed and videotaped tomorrow. [The demo was postponed on 8/7 because of the extended EMU activities.]

The CDR performed the regular daily routine tasks of SOZh life support system maintenance and also the inventory management system (IMS) delta file preparation.

Both crewmembers worked out in their daily 2.5-h program of physical exercise, on the RED expander and TVIS treadmill.

For today’s scheduled troubleshooting of the SAMS (Space Acceleration Measurement System), currently in the MSG awaiting the PFMI experiment initiation, Ed Lu conducted a checkout of its laptop, battery and PCB (power conversion box).  [This isolated the three major components of SAMS in the ICU (Interface Control Unit) drawer.  The results of these activities will determine what spares need to be installed to get SAMS up and running.]

The Russian SKV-1 air conditioner continues to operate OK, as do both KOB thermal loop pumps.  [Background: Since the SKV was deactivated on 8/7 there has been only minimal collection of condensate in the Lab module.  Overall the system has been configured to minimize US segment (USOS) condensate collection during Expedition 7.  SKV was controlling dew point to ~45F (7.2 degC)  and the Lab CCAA (common cabin air assembly) air conditioner was set to control dew point to ~53F (11.7 degC).  Preliminary ground estimates were that ~6 lbs of condensate were absorbed into the atmosphere, and that some additional amounts are unaccounted for in the total amount of condensate that should have been generated over the time the SKV was deactivated.   The crew checked for water on the Airlock (A/L) CCAA heat exchanger jumper hose, to assess whether there was condensation buildup in the A/L and the heat exchanger of the inactive A/L CCAA, one of the local “cold spots” that could attract condensate.]

Initial USSPACECOM notification was received of a conjunction with the second or third stage from a Chinese Long March 4B (CZ-4B) vehicle launched 10/27/02 (object #27551).  Preliminary TCA (time of closest approach): Thursday (8/14), 7:41pm EDT.  The conjunction is of moderate concern due to the relatively small radial miss distance of 172 m.  Cross section: 11.2 squ.m. (120 squ.ft.).  Radar tracking has been increased, and MMC-H continues to monitor.  [The purpose of the CZ-4B is to launch Chinese sun-synchronous orbit meteorological satellites.]

Today’s CEO (crew earth observation) targets, no longer limited in the current LVLH attitude and including the targets of the Lewis & Clark 200-year memorial locations, were Lower Amazon River Basin (pass over the big island in the center of the estuary.  Pointing left and right), Guadalajara, Mexico (nadir pass), Lahore, Pakistan (this capital of Punjab has nearly 4.5 million inhabitants), Berlin, Germany (hot clear weather persists), Denver, Colorado (nadir pass), Pomeroy, Washington (on their return eastward, the Lewis & Clark explorers decided to go overland in order to avoid battling upstream against the strong Snake River, traveling  through the village of Pomeroy, WA), Forks of Clearwater (LEWIS & CLARK SITE:  At Canoe Camp, Oct 1, 1805, one of the diarists reported that most of the men in the party were very weak.  To save them from labor, they made new canoes using the Indian method of burning out the interior of logs rather than using axes), Great Falls, MT (LEWIS & CLARK SITE:  The Missouri River drops more than 400 feet in ten miles here.  The explorers spent more than a month here in June-July 1805, and noted the value of this site for hydro-power generation.  Now the Ryan Dam occupies the site.  The site lies on the river, at the city of Great Falls, where the highway crosses the river, all at nadir), and Mouth of Marias River (LEWIS & CLARK SITE:  The Marias is a large, meandering river that joins the Missouri from the northwest (about 80 miles northeast of Great Falls).  The river valley that includes Lake Elwell Reservoir is at nadir.  On the return trip of the expedition in the summer of 1806, Lewis and nine men temporarily separated from the main party to explore this river and valley).
CEO images can be viewed at the websites

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 1:25pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered Off.  Vozdukh CO2 scrubber is On (in Manual Mode 5/3).  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is operating.  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 26.3; ppO2 (mmHg) — data invalid; ppCO2 (mmHg) — data invalid.
  • SM Transfer Compartment:  Pressure (mmHg) — 741; temperature (deg C) — 20.0.
  • FGB Cabin:  Pressure (mmHg) — 744; temperature (deg C) — 22.3.
  • Node:  Pressure (mmHg) — 735.62; temperature (deg C) — 23.5 (shell); ppO2 (mmHg) — 162.4; ppCO2 (mmHg) — 3.1.
  • U.S. Lab:  Pressure (mmHg) — 737.95; temperature (deg C) — 23.3; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 738.15; temperature (deg C) — 27.0; shell heater temp (deg C) — 23.8, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 24.7
  • PMA-2:  Shell heater temp (deg C) — 16.0.

(n/a = data not available)

Propulsion System (PS):

  • Total propellant load available: 3646 kg (8038 lb) as of 7/31  [SM(774) + FGB(2872) + Progress M(0) +Progress M-1(0)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Autotrack (solar
  • tracking with 45 deg bias for drag reduction).
  • SM batteries:  Battery #5 is off line(Capacity Restoration mode, ROM); battery #6 is in “Cycle” mode; all other batteries (6) are in “Partial Charge” mode.  Batteries #1and #3 showing degraded capacity. 
  • FGB batteries:  Battery #1 is off line (Capacity Restoration mode, ROM); all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-1 is operating; INT-2 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-2 MDM is Off; PL-1 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist).Management.
  • Solar Beta angle:  22.2 deg (magnitude increasing).

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-2 is prime).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SRMS/Canadarm2 based at MBS PDGF #1 with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 11:35am EDT [= epoch]):

  • Mean altitude — 384.3 km
  • Apogee — 389.7 km
  • Perigee — 378.9 km
  • Period — 92.24 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0007989
  • Orbits per 24-hr. day — 15.61
  • Solar Beta Angle — 22.2 deg (magnitude increasing)
  • Mean altitude loss in last 24 hours — 100 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 26982
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.